A newly issued bond pays its coupons once annually. Its coupon rate is 7.7%, its maturity is 20 years, and its yield to maturity is 9.5%.

a. Find the holding-period return for a 1-year investment period if the bond is selling at a yield to maturity of 8.5% by the end of the year. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

holding period return?

b. If you sell the bond after one year, what taxes will you owe if the tax rate on interest income is 40% and the tax rate on capital gains income is 30%? The bond is subject to original-issue discount tax treatment. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

tax on intrest income

tax on capital gain

tot taxes

c. What is the after-tax holding-period return on the bond? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

d. Find the realized compound yield before taxes for a 2-year holding period, assuming that (1) you sell the bond after two years, (2) the bond yield is 8.5% at the end of the second year, and (3) the coupon can be reinvested for one year at a 3% interest rate. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

realized compaund yield before taxes

e. Use the tax rates in (b) above to compute the after-tax 2-year realized compound yield. Remember to take account of OID tax rules. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

after tax 2yr-realized compound yield
Answer #1

Price of the bond can be obtained using the PV function of excel as Price = - PV (Rate, Nper, PMT, FV)

Part (a)

At t = 0, the price of the bond = P0 = - PV (9.5%, 20, 7.7% x 1000, 1000) = 841.38

At t = 1, the price of the bond = P1 = - PV (8.5%, 19, 7.7% x 1000, 1000) = 925.86

Coupon paid in between = C = 7.7% x 1,000 = 77

the holding-period return for a 1-year investment period = (P1 + C) / P0 - 1 = (925.86 + 77) / 841.38 - 1 = 19.19%

Part (b)

If the yield was constant at 9.5%, then

Constant yield price P1* at the end of year 1 = - PV (9.5%, 19, 7.7% x 1000, 1000) = 844.31

Constant yield price P2* at the end of year 2 = - PV (9.5%, 18, 7.7% x 1000, 1000) = 847.52

Implicit interest at the end of year 1 under OID taxation provision = P1* - P0 = 844.31 - 841.38 = $ 2.93

Hence, total interest income, I = Annual coupon + implicit interest = 77 + 2.93 = 79.93

Capital gain, CG = P1 - P1* = 925.86 - 844.31 = 81.55

Tax rate on interest, Ti = 40%; Tax on capital gain, Tcg = 30%

tax on interest income = I x Ti = 79.93 x 40% = 31.97

tax on capital gain = CG x Tcg = 81.55 x 30% = 24.47

total taxes = 31.97 + 24.47 = 56.44

Part (c)

After-tax holding-period return on the bond = (P1 + C - total taxes) / P0 - 1 = (925.86 + 77 - 56.44) / 841.38 - 1 = 12.48%

Part (d)

Price of the bond at the end of year 2 = P2 = - PV (8.5%, 18, 7.7% x 1000, 1000) = 927.56

Coupon reinvestment maturity amount = 77 x (1 + 3%) = 79.31

Hence, 2 years HPR before taxes = (P2 + Coupon reinvestment maturity amount + Coupon at the end of year 2) / P0 - 1 = (927.56 + 79.31 + 77) / 841.38 - 1 = 28.82%

Hence, the realized compound yield before taxes (1 + 28.82%)1/2 - 1 = 13.50%

Part (e)

Taxes in year 2:

Implicit interest in year 2 = P2* - P1* = 847.52 - 844.31 = 3.21

Tax on interest in year 2 = 40% x (Coupon at the end of year 2 + Implicit interest in year 2 + Coupon 1 reinvestment income) = 40% x (77 + 3.21 + 77 x 3%) = 33.01

Capital gain in year 2 = P2 - P2* = 927.56 - 847.52 = 80.04

Capital gain tax for year 2 = 30% x 80.04 = 24.01

Hence, total tax for year 2 = 33.01 + 24.01 = 57.02

Hence, post tax HPR over 2 years = (P2 + Coupon reinvestment maturity amount + Coupon at the end of year 2 - total tax of year 1 - total tax of year 2) / P0 - 1 = (927.56 + 79.31 + 77 - 56.44 - 57.02) / 841.38 - 1 = 17.90%

Hence, the realized compound yield after taxes (1 + 17.90%)1/2 - 1 = 8.58%